Help Please !
A sample of 30 observations is selected from a normal population. The sample mean is 10, and the population standard deviation is 2. Conduct the following test of hypothesis using the 0.10 significance level.
H0: μ ≤ 9
H1: μ > 9
Q1)Is this a one- or two-tailed test?

One-tailed test ?
Two-tailed test ?

Q2)What is the decision rule?

Reject H0 when z > 1.282 ?
Reject H0 when z ≤ 1.282 ?
Q3)What is the value of the test statistic? (Round your answer to 2 decimal places.)
Q4) What is your decision regarding H0?

Reject H0
Fail to reject H0
Q5) What is the p-value? (Round your answer to 4 decimal places.)

Q6) Interpret the p-value? (Round your final answer to 2 decimal places.)
There is a .......%chance of finding a z value this large by ''Sampling Eror'' When H0 is true

Answer by Boreal(15235)

Two tail tests have an equal not equal dichotomy. This is one tail.
The null hypothesis has mean of 9 and the sample has mean 10. The z test will be positive z >1.282
value of test statistic
z=(x-mean)/sigma/sqrt(n)
=(10-9)/2/sqrt(30)=1*sqrt(30)/2=2.74
Reject Ho
p-value is probability z > 2.74, and that is 0.0031.
If this result were by chance, it would occur about 3 times in a 1000. or 0.31% of the time.